OpenStudy (anonymous):

a particle moving along the x axis has the position x(t) at time t with the velocity of the particle given by v(t)=5sin(t^2). At time t=6, the particles position is (4,0). Find the position of the particle when t=7... please and thank you :)

OpenStudy (anonymous):

x(t)=int(5sint^2)=5t/2-5/4sin2t+c 4=5*6/2-5/4sin12+c c=4-15+0.26= -10.74 x(7)=5*7/2-5/4sin14-10.74= =17.5-0.3-10.74=6.46 I am not so sure !